# Physics-informed Neural Mapping and Motion Planning in Unknown Environments

### v1

@@ I Introduction
The arrival time is determined by solving the Eikonal equation, representing the shortest travel time from a starting point to a destination. The gradients of the arrival time guide the generation of a continuous shortest path for motion planning [52, 37]. For example, the Fast Marching Method (FMM) solves the Eikonal equation via grid search over discretized space for path planning [52, 62, 61]. However, these methods do not scale to higher-dimensional robot configuration space. Recent methods solve the Eikonal equation via neural networks in continuous space and scale to higher dimensional motion planning problems [37, 38]. Their results show that the arrival time fields allow path planning much faster than any other motion planner. However, these methods require a known environment and infer arrival time fields through offline training of neural networks. Furthermore, their training process involves the computation of the Laplacian coupled with curriculum learning, which significantly adds to the neural network training times, making them unsuitable for online applications. In contrast, the solution to the mapping problem requires an online methodology to generate the map of an unknown environment.
This paper introduces a novel approach called Active Neural Time Fields (Active NTFields), which quickly maps the arrival time fields in unknown environments in an online manner and uses them for real-time motion planning in complex environments. The main contributions of our proposed work are listed as follows:
@@ II Related Work
Perhaps the closest relevant mapping method is the cost-to-go (c2g) map [9, 57, 8, 21, 30]. This map stores the c2g function in C-Space coordinates, and the gradients of this function guide motion planning paths. Traditionally, the c2g function relies on discrete representations of free space environments, calculated through wavefront [52, 62, 61] or diffusion methods [13, 15, 9, 16]. However, these methods are inefficient and challenging to scale for high DOF conditions. Learning-based approaches have been suggested to learn the c2g function on grid [57, 8] or continuous space [21, 30]. Unfortunately, they necessitate expert c2g functions as supervision data, leading to time-consuming data generation and offline training.
Recently, physics-informed neural networks (PINN) [48] have emerged as promising tools for motion planning. They find path solutions in a fraction of a second and do not require expensive expert trajectories for learning. NTFields [37] and its variant P-NTFields [38] are the first of these methods that directly learn to solve the Eikonal equation without needing expert trajectories. The solution of the Eikonal equation is an arrival time field between the given start and goal. The arrival time field is a type of c2g function whose gradients lead to a path solution. The NTFields were limited by the complex loss landscape of the Eikonal equation. The P-NTFields extended the Eikonal equation via Laplacian and introduced progressive speed scheduling to guide neural network learning. However, these additions significantly increased the training time of P-NTFields, making it unsuitable for online mapping tasks of unknown environments.
@@ III Background
The solution of the above PDE is the shortest arrival time between the given start and goal. The prior methods solved the above equation using the grid-based Fast Marching Method (FMM) to recover the shortest path based on arrival time between the given start and goal [52, 62, 61]. However, these methods lacked continuity and were computationally intractable in higher dimensional spaces. Recent work introduced a Laplacian-based viscosity term in the Eikonal equation and a new factorized formulation of arrival time, which allowed for solving Eq. 1 using neural networks [37, 38]. The factorized arrival time is described as follows:
@@ IV Proposed Method
Figure 1: Active NTFields pipeline: It starts from the camera’s current viewpoint’s perception and samples C-Space points $q_{s},q_{g}$ with ground truth speed fields from existing observations. This online data is used to train the neural networks based on the Eikonal equation loss function, leading to the time field estimation. Finally, the next viewpoint is selected using an exploration strategy, and the local arrival time fields guide the robot to that viewpoint without needing any external motion planner. The process repeats until the given environment is observed and its arrival time fields are reconstructed.
@@ IV-A New Time Field Factorization
It can be seen that we replace $1/{\tau}$ in Equation 2 with ${\log(\tau)^{2}}$. The latter is more flattened and changes less steeply as $\tau$ decreases, preventing sharp features near obstacles. We show in our experiments that the new definition recovers accurate arrival time fields without needing Laplacian or speed scheduling. Furthermore, by solving the Eikonal equation (Eq. 1) via chain rule with the new arrival time definition in Eq. 5, the speed $S$ becomes as follows:
$S(q_{g})=\frac{\tau(q_{s},q_{g})/\log({\tau(q_{s},q_{g})})}{\sqrt{\begin{aligned} \tau(q_{s},q_{g})^{2}\log({\tau(q_{s},q_{g})})^{2}+4{\tau(q_{s},q_{g})}\\ \log({\tau(q_{s},q_{g})})\nabla_{q_{g}}\tau(q_{s},q_{g})\cdot(q_{s}-q_{g})\\ +4{\|q_{s}-q_{g}\|^{2}}\|\nabla_{q_{g}}\tau(q_{s},q_{g})\|^{2}\end{aligned}}}$ (6)
@@ IV-B Neural Model Design
Our neural architecture is inspired by spectral distance (SD) formulation [12, 31]. The SD $d_{w}(q_{s},q_{g})$ between two points $q_{s}$ and $q_{g}$ is an approximate alternative to the shortest arrival time with the following form:
@@
where $\phi_{i}(\cdot)$ and $\lambda_{i}$ are the eigenfunctions and eigenvalues, respectively, of the Laplace operator $\Delta$ in the given collision-free space and $w(\cdot)$ is a weight function. The eigenfunction of the Laplace operator in the given collision-free space implies $-\Delta_{q}\phi_{i}(q)=\lambda_{i}\phi_{i}(q)$ when $q\in\mathcal{Q}_{free}$, whereas $\nabla_{\mathbf{n}}\phi_{i}(q)=0$ when $q\in\partial\mathcal{Q}_{free}$. In practice, $k$ smallest eigenvalues with corresponding eigenfunctions are used to estimate the spectral distance.
By interpreting eigenfunctions as a latent feature space, the SD formulation demonstrates that the difference in feature space vectors between the start and goal points sufficiently represents the shortest arrival time. Our neural time field network design is based on this SD formulation. Additionally, to enhance the representation capability, we integrate random Fourier positional encoding [58] with the SIREN neural network [53].
@@
The above squared subtraction is also relevant to the symmetric operator introduced in NTFields [37]. This operator enforces the arrival time field’s symmetric property, i.e., the arrival time from start to goal and from goal to start should be the same. Specifically, let latent encodings of start and goal configurations be denoted as $a$ and $b$. The symmetric operator in [37] was defined as the concatenation of min and max, i.e., $[\max(a,b),\min(a,b)]$. The other choice of symmetric operators could be subtraction, addition, or averaging. Hence, the squared subtraction operator in Eq. 9, inspired by spectral distance, also acts as a symmetric operator, ensuring the symmetric property of the arrival time field. Furthermore, we observed that the squared subtraction operator instead of concatenation of min and max reduces the feature size, leading to neural network training efficiency. Finally, the resulting features from Equation 9 are given to another multi-layer perceptron, denoted as $g$, that predicts the $\tau$. In summary, our neural time field network with parameters $\theta$ is summarized as follows:
@@ IV-C Speed Inference
The neural framework mentioned above predicts $\tau(q_{s},q_{g})$ for the given start and goal pairs. We use the predicted $\tau$ and its gradients with respect to start $\nabla_{q_{s}}\tau(q_{s},q_{g})$ and goal $\nabla_{q_{g}}\tau(q_{s},q_{g})$ to estimate the speeds $S(q_{s})$ and $S(q_{g})$ according to Eq. 6. These predicted speeds at the start and goal are then utilized for training the network by comparing them against the expert speed model, as described in section V-G.
@@ IV-D Path Inference
Once the neural time field network is trained, we utilize its predictions and gradients to recover the path solution between any start and goal as follows. First, the network predicts $\tau(q_{s},q_{g})$, and we use it to predict speeds (Section V-C) and also to parameterize Eq. 5 to determine the arrival time field $T(q_{s},q_{g})$. Next, we recover the path solution in the same way as in NTFields [37], i.e., we follow gradients bidirectionally from start to goal and from goal to start as described below:
@@
The $\alpha$ is a user-defined step-size scaling factor. The above iterative procedure is repeated until the waypoints from the start and the goal are within the threshold. It should be emphasized that gradient descent does not ensure convergence, and the algorithm may terminate unsuccessfully after a predefined number of iterations.
@@ IV-E Active Neural Time Fields
This section describes our approach to gathering data online in unknown environments for training our neural network on the fly. The prior work assumes the environment to be known, and therefore, the dataset for training the neural time field models is gathered offline. The dataset comprises randomly sampled start and goal points and their ground truth speed values. As described in Section IV (Eq. 4), the ground truth speed $S^{*}$ is computed for each sampled point based on its minimum distance from the obstacles. The NTFields [37] and P-NTFields [38] can directly acquire this distance from a known environment by calculating the distance between the point and the nearest obstacle mesh. However, in our setting, the environment is unknown, and the robot needs to explore, obtain data, and determine its ground truth speed values to train the neural networks. Therefore, this section provides procedures to actively create such data and define their ground truth speed for online training of our neural networks.
@@ IV-E1 Local Perception Processing
Next, we need to determine the ground truth speed values for all sampled points. In our setting, the actual minimum distance to the nearest obstacle is unknown since the environment will only be partially observed during exploration. This contrasts with prior methods where the environment mesh is provided, and the nearest distance to obstacles is readily available. Therefore, in our framework, we revert to approximating the minimum distance of sampled points to the obstacles. We calculate the distance between the point and the nearest surface point in the given scan. Although using all lidar points or depth points can be more accurate for distance estimation, it is computationally expensive. Therefore, we only use $N$ sampled rays to select the nearest obstacle surface. For 3D point robot, workspace sampled points $\boldsymbol{\mathrm{x}}_{\{n\}}$ are the C-Space points $\boldsymbol{\mathrm{q}}_{\{n\}}$. However, in the case of a higher-dimension manipulator, we need a further operation to get C-Space samples and their expert speed values. Specifically, we use $\boldsymbol{\mathrm{x}}_{\{n\}}$ as the manipulator end effect positions and determine their configuration samples $\boldsymbol{\mathrm{q}}_{\{n\}}$ via inverse kinematic solution. Next, for these C-Space samples, we determine the robot surface points to compute the minimum distance to the obstacle. Once the approximate minimum distance of all sampled points from the obstacle is available, we define the expert speed model based on Eq. 4 (Section IV). In summary, this module processes local perception, creating sample points $\boldsymbol{\mathrm{q}}_{\{n\}}$ and their expert speed values $S^{*}_{\{n\}}$.
@@ IV-E2 Online Training
We pass the sampled start and goal pairs through our neural time field network to obtain the $\tau$. The $\tau$ and its gradients parametrize Eq. 6 to infer the speed $S$ (Section V-C). Next, we compute the loss of predicted speed against ground truth speed as follows and use it to train our neural networks for a fixed number of iterations.
@@
FMMOursP-NTFieldsNTFields
Figure 3: We compare our method with NTFields, P-NTFields, and FMM on a maze environment for time field generation. Our method recovers the correct result within 15 seconds, whereas P-NTFields and NTFields take 18 and 10 minutes, respectively.
@@ IV-E3 Next Viewpoint Selection
To determine the next best view for exploration, we randomly sample collision-free points from the frontier of the explored occupancy map, considering these as candidates for the next best view. We then compute the arrival time to these points using our neural model from the robot’s current location and select the point with the minimal arrival time. The path from the robot’s current location to this next best view is calculated using the arrival time field, following the procedure described in Section IV-D.
@@ V-A Training Efficiency Analysis
This section demonstrates our method performance over NTFields [37] and P-NTFields [38]. Fig. 3 shows the speed and time fields of our method, NTFields, and P-NTFields in comparison to FMM as ground truth. The color shows the speed fields and the contours show the arrival time fields from a start point. From Fig. 3 contours, our approach and P-NTFields get a similar result as FMM, while NTFields generate incorrect time fields shown in the red box. To quantitatively compare the results of all PINN methods, we sample points from their predicted arrival time and compute absolute error to the ground truth time field value given by the FMM. Our time fields have an error of $0.044\pm 0.050$, P-NTFields have an error of $0.048\pm 0.047$, and NTFields have an error of $0.44\pm 0.42$. Note that the NTField error is almost ten times higher. Regarding training times, our method takes 15 seconds, while P-NTFields and NTFields take 18 minutes and 10 minutes, respectively. While our method and P-NTFields have similar results, the latter is about 72 times faster than the former in terms of training speed. Hence, this performance comparison validates our method’s effectiveness and its suitability for online continual learning in mapping tasks.

### v2

@@ I Introduction
The arrival time is determined by solving the Eikonal equation, as described in detail in Section III. The arrival time represents the shortest travel time from a starting point to a destination. The gradients of the arrival time guide the generation of a continuous shortest path for motion planning [56, 41]. For example, the Fast Marching Method (FMM) solves the Eikonal equation via grid search over discretized space for path planning [56, 67, 66]. However, these methods do not scale to higher-dimensional robot configuration space. Recent methods solve the Eikonal equation via neural networks in continuous space and scale to higher dimensional motion planning problems [41, 40, 42, 57, 33]. Their results show that the arrival time fields allow path planning much faster than any other motion planner. However, these methods require a known environment and infer arrival time fields through the offline training of neural networks. Additionally, their process to solve the Eikonal equation is either slow or prone to errors in complex environments, making them unsuitable for online applications. For more details on these previous neural Eikonal equation solvers, please refer to Section III. Experimental comparisons with previous methods are provided in Section V.
In contrast to these earlier methods that require a known environment and a computationally intensive training process, the solution to the mapping problem demands a fast, efficient, and reliable methodology for generating a map of an unknown environment. Therefore, the objective of this work is to solve the Eikonal equation using a neural network in unknown environments based on local observations without relying on costly training strategies and complex loss functions.
To achieve the above-mentioned objectives, this paper introduces a novel approach called Active Neural Time Fields (Active NTFields), which quickly maps the arrival time fields in unknown environments in an online manner and uses them for real-time motion planning in complex environments. The main contributions of our proposed work are listed as follows:
@@ II Related Work
Perhaps the closest relevant mapping method is the cost-to-go map [10, 62, 9, 22, 32]. This map stores the cost-to-go function in C-Space coordinates, and the gradients of this function guide motion planning paths. Traditionally, the cost-to-go function relies on discrete representations of free space environments, calculated through wavefront [56, 67, 66] or diffusion methods [14, 16, 10, 17]. However, these methods are inefficient and challenging to scale for high DOF conditions. Learning-based approaches have been suggested to learn the cost-to-go function on grid [62, 9] or continuous space [22, 32]. Unfortunately, they necessitate expert cost-to-go functions as supervision data, leading to time-consuming data generation and offline training.
Recently, physics-informed neural networks (PINN) [52] have emerged as promising tools for motion planning. They find path solutions in a fraction of a second and do not require expensive expert trajectories for learning. NTFields [41] and its variant P-NTFields [40] are the first of these methods that directly learn to solve the Eikonal equation without needing expert trajectories. The solution of the Eikonal equation is an arrival time field between the given start and goal. The arrival time field is a type of cost-to-go function whose gradients lead to a path solution. The NTFields were limited by the complex loss landscape of the Eikonal equation. The P-NTFields extended the Eikonal equation via Laplacian and introduced progressive speed scheduling to guide neural network learning. However, these additions significantly increased the training time of P-NTFields, making it unsuitable for online mapping tasks of unknown environments.
@@ III Background
The solution of the above PDE is the shortest arrival time between the given start and goal. The prior methods solved the above equation using the grid-based FMM to recover the shortest path based on arrival time between the given start and goal [56, 67, 66]. However, these methods lacked continuity and were computationally intractable in higher dimensional spaces. Recent work introduced a Laplacian-based viscosity term in the Eikonal equation and a new factorized formulation of arrival time, which allowed for solving Eq. 1 using neural networks [41, 40]. The factorized arrival time is described as follows:
@@ IV Proposed Method
Figure 1: The process begins by capturing the scene from the camera’s current viewpoint and sampling C-space points $q_{s}$ and $q_{g}$. Each sampled point pair $(q_{s},q_{g})$ is then input into the neural network, as shown in Fig. 2, to predict the corresponding arrival times $T$. The predicted speed field is derived using the Eikonal equation, as detailed in Eq. 6. The neural network is trained based on the loss between the ground truth speed and the predicted speed, as outlined in Eq. 12. Finally, the next viewpoint is selected using an exploration strategy, and the local arrival time fields guide the robot to that viewpoint without requiring any external motion planner. This process repeats until the entire environment is observed and its arrival time fields are reconstructed.
@@ IV-A New Time Field Factorization
It can be seen that we replace $1/{\tau}$ in Eq. 2 with ${\log(\tau)^{2}}$. The latter is more flattened and changes less steeply as $\tau$ decreases, preventing sharp features near obstacles. We show in our experiments that the new definition recovers accurate arrival time fields without needing Laplacian or speed scheduling. Furthermore, by solving the Eikonal equation (Eq. 1) via chain rule with the new arrival time definition in Eq. 5, the speed $S$ becomes as follows:
${S(q_{g})=1/\|\nabla_{q_{g}}(\log(\tau(q_{s},q_{g}))^{2}\|q_{s}-q_{g}\|)\|}$ (6)
@@ IV-B Neural Model Design
Our neural architecture is inspired by spectral distance (SD) formulation [13, 34]. The SD $d_{w}(q_{s},q_{g})$ between two points $q_{s}$ and $q_{g}$ is an approximate alternative to an arrival time with the following form:
@@
where $\phi_{i}(\cdot)$ and $\lambda_{i}$ are the eigenfunctions and eigenvalues, respectively, of the Laplace operator $\Delta$ in the given collision-free space and $w(\cdot)$ is a weight function. The eigenfunction of the Laplace operator in the given collision-free space implies $-\Delta_{q}\phi_{i}(q)=\lambda_{i}\phi_{i}(q)$ when $q\in\mathcal{Q}_{free}$, whereas $\nabla_{\mathbf{n}}\phi_{i}(q)=0$ when $q\in\partial\mathcal{Q}_{free}$ and $\nabla_{\mathbf{n}}$ denotes the directional derivative along the outward normal of the collision-free space boundary.
In practice, the spectral distance is computed using the $k$ smallest eigenvalues and their corresponding eigenfunctions. Although SD does not produce the exact shortest arrival time, it provides a robust global approximation sufficient to retrieve a path solution, as illustrated in Fig. 3. This attribute of SD to capture global structure inspires our neural architecture design to capture the global structure of the time field.
This SD formulation inspires our neural time field network to approximate the shortest arrival time. To further improve its representational capacity, we integrate random Fourier positional encoding [63] with the SIREN neural network architecture [58].
@@
The above squared subtraction is also relevant to the symmetric operator introduced in NTFields [41]. This operator enforces the arrival time field’s symmetric property, i.e., the arrival time from start to goal and from goal to start should be the same. Specifically, let latent encodings of start and goal configurations be denoted as $a$ and $b$. The symmetric operator in [41] was defined as the concatenation of min and max, i.e., $[\max(a,b),\min(a,b)]$. The other choice of symmetric operators could be subtraction, addition, or averaging. Hence, the squared subtraction operator in Eq. 9, inspired by spectral distance, also acts as a symmetric operator, ensuring the symmetric property of the arrival time field. Furthermore, we observed that the squared subtraction operator instead of concatenation of min and max reduces the feature size, leading to neural network training efficiency. Finally, the resulting features from Eq. 9 are given to another multi-layer perceptron, denoted as $g$, that predicts the $\tau$. In summary, our neural time field network with parameters $\theta$ is summarized as follows:
@@ IV-C Speed Inference
The neural framework mentioned above predicts $\tau(q_{s},q_{g})$ for the given start and goal pairs. We use the predicted $\tau$ and its gradients with respect to start $\nabla_{q_{s}}\tau(q_{s},q_{g})$ and goal $\nabla_{q_{g}}\tau(q_{s},q_{g})$ to estimate the speeds $S(q_{s})$ and $S(q_{g})$ according to Eq. 6. These predicted speeds at the start and goal are then utilized to train the network by comparing them against the expert speed model, as described in Section III.
@@ IV-D Path Inference
Once the neural time field network is trained, we utilize its predictions and gradients to recover the path solution between any start and goal as follows. First, the network predicts $\tau(q_{s},q_{g})$, and we use it to predict speeds (Section IV-A) and also to parameterize Eq. 5 to determine the arrival time field $T(q_{s},q_{g})$. Next, we recover the path solution in the same way as in NTFields [41], i.e., we follow gradients bidirectionally from start to goal and from goal to start as described below:
@@
The $\alpha$ is a user-defined step-size scaling factor. The above iterative procedure is repeated until the waypoints from the start and the goal are within the threshold. It should be emphasized that gradient descent does not ensure convergence, and the algorithm may terminate unsuccessfully after a predefined number of iterations. To address this issue, the learned arrival time field maps can also be used more robustly by incorporating them into search [24] or optimization [69, 4] techniques. These methods will guided by the learned arrival time field map, yielding effective path solutions fast.
@@ IV-E Active Neural Time Fields
This section describes our approach to gathering data online in unknown environments for training our neural network on the fly. The prior work assumes the environment to be known, and therefore, the dataset for training the neural time field models is gathered offline. The dataset comprises randomly sampled start and goal points and their ground truth speed values. As described in Section IV-A (Eq. 4), the ground truth speed $S^{*}$ is computed for each sampled point based on its minimum distance from the obstacles. The NTFields [41] and P-NTFields [40] can directly acquire this distance from a known environment by calculating the distance between the point and the nearest obstacle mesh. However, in our setting, the environment is unknown, and the robot needs to explore, obtain data, and determine its ground truth speed values to train the neural networks. Therefore, this section provides procedures to actively create such data and define their ground truth speed for online training of our neural networks.
@@ IV-E1 Local Perception Processing
Next, we need to determine the ground truth speed values for all sampled points. In our setting, the actual minimum distance to the nearest obstacle is unknown since the environment will only be partially observed during exploration. This contrasts with prior methods where the environment mesh is provided, and the nearest distance to obstacles is readily available. Therefore, in our framework, we revert to approximating the minimum distance of sampled points to the obstacles. We calculate the distance between the point and the nearest surface point in the given scan. Although using all lidar points or depth points can be more accurate for distance estimation, it is computationally expensive. Therefore, we only use $N$ sampled rays to select the nearest obstacle surface. For 3D point robot, workspace sampled points $\boldsymbol{\mathrm{x}}_{\{n\}}$ are the C-Space points $\boldsymbol{\mathrm{q}}_{\{n\}}$. However, in the case of a higher-dimension manipulator, we need a further operation to get C-Space samples and their expert speed values. Specifically, we use $\boldsymbol{\mathrm{x}}_{\{n\}}$ as the manipulator end effect positions and determine their configuration samples $\boldsymbol{\mathrm{q}}_{\{n\}}$ via inverse kinematic solution. Next, for these C-Space samples, we determine the robot surface points to compute the minimum distance to the obstacle. Once the approximate minimum distance of all sampled points from the obstacle is available, we define the expert speed model based on Eq. 4 (Section IV-A). In summary, this module processes local perception, creating sample points $\boldsymbol{\mathrm{q}}_{\{n\}}$ and their expert speed values $S^{*}_{\{n\}}$.
@@ IV-E2 Online Training
We pass the sampled start and goal pairs through our neural time field network to obtain the $\tau$. The $\tau$ and its gradients parametrize Eq. 6 to infer the speed $S$ (Section IV-A). Next, we compute the loss of predicted speed against ground truth speed and use it to train our neural networks for a fixed number of iterations.
@@
FMMOursP-NTFieldsNTFieldsSD
Figure 3: We compare our method with NTFields, P-NTFields, FMM, and SD on a maze environment for time field generation. Our method recovers the correct result within 15 seconds, whereas P-NTFields and NTFields take 18 and 10 minutes, respectively. Although SD does not recover the exact time field, it provides a smooth global approximation, which inspired aspects of our architecture design.
@@ IV-E3 Next Viewpoint Selection
To determine the next best view for exploration, we randomly sample collision-free points from the frontier of the explored occupancy map, considering these as candidates for the next best view. We then compute the arrival time to these points using our neural model from the robot’s current location and select the point with the minimal arrival time. The path from the robot’s current location to this next best view is calculated using the arrival time field, following the procedure described in Section IV-E.
@@ V-A Training Efficiency Analysis
This section demonstrates our method performance over NTFields [41] and P-NTFields [40]. Fig. 3 shows the speed and time fields of our method, NTFields, P-NTFields, and SD in comparison to FMM as ground truth. The color shows the speed fields and the contours show the arrival time fields from a start point. From Fig. 3 contours, our approach and P-NTFields get a similar result as FMM, while NTFields generate incorrect time fields. Although the SD contours are not perfectly aligned with those of FMM, they offer a smooth and globally consistent approximation of the arrival time fields, which inspired elements of our feature design for learning the shortest arrival times.
To quantitatively compare the results of all PINN methods, we sample points from their predicted arrival time and compute absolute error to the ground truth time field value given by the FMM. Our time fields have an error of $0.044\pm 0.050$, P-NTFields have an error of $0.048\pm 0.047$, and NTFields have an error of $0.44\pm 0.42$. Note that the NTField error is almost ten times higher. Regarding training times, our method takes 15 seconds, while P-NTFields and NTFields take 18 minutes and 10 minutes, respectively. While our method and P-NTFields have similar results, the latter is about 72 times faster than the former in terms of training speed. Hence, this performance comparison validates our method’s effectiveness and its suitability for online continual learning in mapping tasks.